OpenStudy (anonymous):

PLEASE HELP I REALLY NEED TO PASS! Find The Product. (t-sa)(9t+6sa)

Mehek (mehek14):

use FOIL

OpenStudy (calculusxy):



OpenStudy (calculusxy):

I am not sure if this is correct ...

OpenStudy (calculusxy):

sorry. one moment

Mehek (mehek14):

\(\tt{(t-sa)(9t+6sa)\\t*9t=9t^2\\-sa*9t=-9tsa\\t*6sa=6tsa\\-sa*6sa=-6sa^2\\9t^2-9tsa+6tsa-6sa^2}\)

Mehek (mehek14):

you have to combine like terms

Mehek (mehek14):

\(\tt{-9tsa+6tsa=-3tsa}\)

OpenStudy (calculusxy):

yes that's correct

OpenStudy (freckles):